### v1

# Revising Multimodal VAEs with Diffusion Decoders

Abstract
Multimodal VAEs often struggle with generating high-quality outputs, a challenge that extends beyond the inherent limitations of the VAE framework. The core issue lies in the restricted joint representation of the latent space, particularly when complex modalities like images are involved. Feedforward decoders, commonly used for these intricate modalities, inadvertently constrain the joint latent space, leading to a degradation in the quality of the other modalities as well. Although recent studies have shown improvement by introducing modality-specific representations, the issue remains significant. In this work, we demonstrate that incorporating a flexible diffusion decoder specifically for the image modality not only enhances the generation quality of the images but also positively impacts the performance of the other modalities that rely on feedforward decoders. This approach addresses the limitations imposed by conventional joint representations and opens up new possibilities for improving multimodal generation tasks using the multimodal VAE framework. Our model provides state-of-the-art results compared to other multimodal VAEs in different datasets with higher coherence and superior quality in the generated modalities
1 Introduction
Deep learning has revolutionized many fields through its ability to extract and learn patterns that are otherwise difficult through other means (LeCun et al., 2015; Schmidhuber, 2015). One area that has seen immense progress is representation learning, where deep neural networks learn rich representations that disentangle the underlying explanatory factors of data by projecting the data into a latent space (Bengio et al., 2013). Variational autoencoders (VAEs) have emerged as a powerful framework for learning interpretable representations in an unsupervised manner (Kingma and Welling, 2014; Rezende et al., 2014). VAEs were initially built to process a single modality, such as an image. However, in the real world, data often comes from multiple modalities that are interconnected and expressed in different formats (Baltrusaitis et al., 2019). Multimodal learning aims to build models that can process and relate information from various modalities, such as vision, language, and other modalities together (Ngiam et al., 2011). Incorporating this multimodal information can have a huge impact on the performance of the model on tasks that require understanding from multiple perspectives (Srivastava and Salakhutdinov, 2014).
Multimodal variational autoencoders extend the VAE framework to learn joint representations across modalities (Ngiam et al., 2011; Suzuki et al., 2017; Wu and Goodman, 2018). By capturing the correlations between modalities, these models can learn robust and informative representations. These learned multimodal representations can be used in different ways, including transfer learning in downstream tasks (Hwang et al., 2021; Jing et al., 2024), and they also provide a framework to perform unconditional and conditional generation across the different modalities (Shi et al., 2019; Wesego and Rooshenas, 2023).
Figure 1: Architecture of our model where modality-specific encoders map the input data from different modalities into latent space which are then fused using PoE and passed to the respective decoders.
Figure 2: Unconditional sampling technique that leverages a score-based model to transform a Gaussian noise distribution into the PoE posterior distribution.
Despite their promising potential, multimodal VAEs continue to face significant challenges in effectively integrating diverse modalities, generating high-quality samples, and scaling to high-dimensional data (Daunhawer et al., 2022; Wesego and Rooshenas, 2023). Various prior works have attempted to address these issues by proposing alternative ELBO objectives, yet the core problems still persist in multimodal VAEs (Sutter et al., 2021, 2020). Due to these ongoing limitations, most multimodal studies have been constrained to low-dimensional datasets such as MNIST or PolyMNIST (LeCun and Cortes, 2010; Sutter et al., 2020). Additionally, Daunhawer et al. (2022) highlights the generative discrepancies in mixture-based models and raises concerns about the practicality of these models in real-world applications.
On the other hand, diffusion models have emerged as a powerful class of generative models, gaining substantial attention in recent years for their ability to produce high-quality samples Ho et al. (2020); Dhariwal and Nichol (2021). These models operate by gradually transforming a data distribution into a noise distribution, with the goal of learning to reverse this process. The forward process is a Markovian sequence that incrementally adds noise to the data distribution over several timesteps, while the reverse process, typically modeled by a neural network, learns to remove the noise and reconstruct the data (Ho et al., 2020). A notable advantage of diffusion models is their capacity to generate exceptionally high-quality samples, achieving state-of-the-art performance in various generative tasks, including image synthesis and inpainting (Saharia et al., 2022; Lugmayr et al., 2022). Preechakul et al. (2022) introduced diffusion autoencoder, where the encoder resembles that of a VAE, but the decoder is modeled using a diffusion process. The key difference between a diffusion decoder and a traditional feedforward decoder used in VAEs is that the diffusion decoder conditions the generation process on the latent representation rather than directly feeding it as input. This approach enables the diffusion autoencoder to learn meaningful representations without being constrained by the output size.
In this study, we propose a novel multimodal generative model that leverages the strengths of both diffusion autoencoders and multimodal VAEs by learning a unified joint representation across multiple modalities. In traditional multimodal VAEs, joint representation often becomes restrictive, especially when rich representations are required for complex modalities like images, leading to the need for modality-specific representations Palumbo et al. (2023); Hwang et al. (2021). However, we demonstrate that by employing a flexible diffusion decoder, the joint representation alone suffices to generate high-quality outputs not only for complex modalities but also for other modalities using feedforward decoders.
To achieve this, we use diffusion decoders for modalities where feedforward decoders have shown limitations, such as images, ensuring high-quality sample generation and improved representation capacity. Conversely, for modalities where feedforward or transformer-based decoders have proven effective, like text or sparse outputs, we retain standard VAE decoders. This hybrid approach allows our model to harness the strengths of each generative model class while effectively addressing their respective limitations.
Our model achieves better coherence in conditional generation among the generated modalities while generating high-quality samples. We use the product of experts (PoE) to fuse the representations from different modalities (Wu and Goodman, 2018). This approach enables our model to capture the correlations and dependencies between modalities. For unconditional generation tasks, where the goal is to generate samples from all modalities simultaneously without conditioning on any specific modality, we propose the use of an auxiliary model. This auxiliary model transforms a Gaussian noise distribution into the approximate posterior product of experts distribution. By leveraging this technique, our model gains increased flexibility and can perform well in both conditional and unconditional generation settings.
Figure 2 illustrates the general design and architecture of our proposed model. The modality-specific encoders map the input data from different modalities into latent representations, which are then fused using the PoE mechanism. The fused representation is then passed to the respective decoders, which can be either diffusion autoencoders or standard VAEs, depending on the modality. Figure 2 shows how to perform unconditional sampling where the score-based model is employed to provide the necessary input distributions to the decoders.
Our proposed model demonstrates superior performance compared to previous variants of multimodal variational autoencoder models. Our key contributions can be summarized as follows:
  • We introduce a novel generative multimodal autoencoder architecture that seamlessly integrates diffusion autoencoders with standard autoencoders. This hybrid design allows our model to generate high-quality samples across modalities without compromising the coherence and consistency among the generated modalities. By doing so, our architecture strikes an optimal balance between sample quality and cross-modal coherence.
  • To enhance the model’s performance in unconditional generation settings, we propose a novel sampling technique that employs an auxiliary score-based model that transforms a Gaussian noise distribution into the approximate distribution of the PoE posterior to improve the model’s ability to generate coherent and diverse samples across all modalities in an unconditional manner.
  • Our approach can be scaled to high-dimensional data modalities with good generative quality compared to previous approaches
Through these contributions, our model addresses the limitations of previous multimodal VAE approaches, while retaining their desirable properties and extending their capabilities to handle high-dimensional data modalities effectively. Our approach paves the way for more powerful and flexible multimodal generative models capable of generating high-quality samples with enhanced coherence.
2 Related Works
As the path to increase the performance and understanding of neural networks continues, multimodal learning has emerged as a promising research direction in the deep learning field that has started from early works such as Ngiam et al. (2011); Tsai et al. (2019). Ngiam et al. (2011) demonstrated the potential of multimodal deep learning models in leveraging information from multiple modalities to improve performance on multiple tasks by learning combined features. As variational autoencoders (VAEs) gained traction for simultaneously learning representations and generation models within individual modalities, researchers soon explored extending the VAE framework to handle multiple modalities simultaneously, giving rise to multiple variants of multimodal VAEs.
Suzuki et al. (2017) first presented a model that is trained jointly and can accept multiple modalities. They propose training additional encoders for each modality and minimizing the KL loss between the joint ones to handle missing modalities. Wu and Goodman (2018) proposed a product of experts approach to encode the modalities jointly. They posterior is constructed as $q_{\phi}(\mathbf{z|x})=p(\mathbf{z})\prod_{i}^{M}q_{\phi}(\mathbf{z}|\mathbf{x_{i}})$. This approach allowed different formulations of the ELBO in future works because the product of experts can be calculated in a closed form, and it allows the generation of missing modalities by simply plugging $1$ in the product for the missing modalities. The model proposed by Wu and Goodman (2018), which is called MVAE, generates images comparative to a unimodal VAE but has low coherence amongst the generated modalities.
Shi et al. (2019) proposed a mixture of experts posterior formulation that samples one modality at a time from the set containing all the modalities. Their model shows better coherence but, in general, has degraded quality, especially visible in image modalities. Sutter et al. (2021) further expanded the subsets in the mixture by proposing a mixture of product of experts posterior. Daunhawer et al. (2022) showed that models that subsample will have a loss in generative quality because of reconstructing all available modalities from a few of them. Other approaches by Lee and Pavlovic (2021); Palumbo et al. (2023); Sutter et al. (2020) explored splitting the latent space into modality-specific and shared representations to improve the generative quality of multimodal VAEs.
In the work by Hwang et al. (2021), the authors propose an evidence lower bound (ELBO) objective derived from an information-theoretic perspective that uses a product of experts approach. Their main objective is to learn a representation that minimizes the conditional total correlation between the joint conditional and the factorized conditionals. Their final objective is a convex combination of two terms: one based on the variational information bottleneck (VIB) and the other based on the conditional VIB.
Alternatively, Wesego and Rooshenas (2023); Bounoua et al. (2024) proposed to use unimodal VAEs to represent each modality and use an additional neural network to fuse the modalities instead of a product/mixture of experts using a two-step training process. They both select a score-based network to join the modalities in the latent space and sample from the joint distribution. Rombach et al. (2021) proposed text-to-image generation that employs a high-quality autoencoder to compress images into latent space, allowing for diffusion processes to occur in this reduced dimensionality. They integrated CLIP models using cross-attention to effectively represent and condition text inputs. Building upon this,  Xu et al. (2023); Tang et al. (2023) proposed multi-modal latent diffusion models that can generate any modality from any given condition. They expand Rombach et al. (2021) work to use multiple diffusion flows to achieve this. It’s worth noting that these advanced models require training on extremely large-scale datasets which requires extensive computational resources for the training processes.
Preechakul et al. (2022) introduced diffusion autoencoder as a means for diffusion models to generate images and learn representations. They transformed the diffusion model into an autoencoder form where an encoder learns a latent space that will be used in the reconstruction process.  Hudson et al. (2023) modified the diffusion autoencoder to reconstruct multiple views in the diffusion decoder.  Wang et al. (2023) proposed a diffusion autoencoder architecture where the latent variable $\mathbf{z}$ is encouraged to learn disentangled meaningful representation by regularizing the objective with mutual information among $\mathbf{x}$ and $\mathbf{z}$.
In this work, we focus on improving multimodal VAEs by proposing an objective that flexibly combines both diffusion autoencoder and standard VAEs. Overall, Daunhawer et al. (2022) illustrated that the current multimodal VAEs are very limited due to their lack of generation capability as the modalities become more complex. This is caused by how the multimodal ELBO objective is constructed and by the limitation in VAEs to generate high-quality images (Pandey et al., 2022). Our approach enhances the generative quality of multimodal VAEs while still preserving the fundamental framework and strengths of the multimodal VAE architecture in conditional and unconditional settings.
3 Methodology
We first present our method starting with diffusion autoencoders. Subsequently, we introduce the multimodal VAE framework that incorporates diffusion autoencoders. The section concludes by explaining how to perform both conditional and unconditional sampling techniques.
3.1 Diffusion Autoencoder
A diffusion model can be formulated as a latent variable model where the marginal distribution is expressed as $p_{\theta}(\mathbf{x}_{0})=\int p_{\theta}(\mathbf{x}_{0:T})\mathrm{d}\mathbf{x}_{1:T}$ (Ho et al., 2020). The combined probability distribution $p_{\theta}(\mathbf{x}_{0:T})=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is the reverse joint process where $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{0},\mathbf{I})$. One can sample any intermediate time step $t$ in the forward process starting from the data point $\mathbf{x}_{0}$ using the Gaussian distribution $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ to add noise to the data. A neural network is trained to reverse this process where $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t))$. Ho et al. (2020) observed that using a simplified objective and predicting the noise is better and easier to implement. The objective is shown as follows in equation  1:
$L(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\epsilon_{\theta}(\mathbf{x}_{t},t)\right\|^{2}\right]$ (1)
where $\mathbf{x_{t}}=\sqrt{\bar{\alpha}_{t}}\mathbf{x_{0}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$ and $\bar{\alpha}_{t}$ are schedule hyperparameters.
A diffusion autoencoder combines the architectural setup of an autoencoder with a diffusion decoder (Preechakul et al., 2022). An encoder $q(\mathbf{z|x})$ projects the data into a latent representation $\mathbf{z}$. This latent representation is fed to the diffusion decoder which generates samples conditioned on $\mathbf{z}$. By training the model this way, the diffusion model can be used as a powerful representation where the encoder learns a meaningful latent representation of the data similar to an autoencoder (Hudson et al., 2023). The objective for a diffusion autoencoder is very similar to the diffusion objective, except we have additional conditioning coming from the encoder. It is trained by minimizing $\mathbb{E}_{\mathbf{x_{0},\epsilon_{t}}}[\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{z},t)\right\|^{2}]$ where $\mathbf{z}$ comes from the encoder network $q(\mathbf{z|x})$.
3.2 Multimodal VAE with Diffusion Decoder
Assuming that we have M modalities, the observed data is represented as $\mathbf{x}={\mathbf{x}^{(1)},\mathbf{x}^{(2)},\ldots,\mathbf{x}^{(M)}}$. A multimodal ELBO objective optimizes the lower bound of the log-likelihood $\log p_{\theta}(\mathbf{x})$. In the multimodal setting, we can write the ELBO as shown in equation  2 where $\mathbf{x}$ represents the multimodal data.
$\begin{split}\log p(\mathbf{x})\geq\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]-D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})\|p(\mathbf{z}))\end{split}$ (2)
Daunhawer et al. (2022) showed that for some subset $\mathcal{S}$ from the combination of the modalities, a lower bound can be derived, shown in equation 3, where different subsets can give rise to different multimodal VAE objectives of previous works of MoPoE, MMVAE, and MVAE Sutter et al. (2021); Shi et al. (2019); Wu and Goodman (2018).
$\begin{split}\log p(\mathbf{x})\geq&\sum_{A\in\mathcal{S}}\omega_{A}\Big{\{}\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x}_{A})}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]-D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x}_{A})\|p(\mathbf{z}))\Big{\}}.\end{split}$ (3)
If we assume the modalities are conditionally independent given the latent representation $\mathbf{z}$, we can express it as $p_{\theta}(\mathbf{x_{1}},\mathbf{x_{2}},...,\mathbf{x_{M}})=p({\mathbf{z}})p_{\theta}(\mathbf{x}_{1}|\mathbf{z})p_{\theta}(\mathbf{x}_{2}|\mathbf{z})...p_{\theta}(\mathbf{x}_{M}|\mathbf{z})$. Then, the ELBO can be written as shown in equation 4:
$\begin{split}\log p(\mathbf{x})\geq&\sum_{A\in\mathcal{S}}\omega_{A}\Bigg{\{}\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x}_{A})}\left[\sum_{i}^{M}\log p_{\theta}(\mathbf{x}_{i}|\mathbf{z})\right]-D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x}_{A})\|p(\mathbf{z}))\Bigg{\}}.\end{split}$ (4)
In our case, we want to use diffusion decoders for image modalities as the generative model while we still want to maintain support for standard VAEs for some modalities. In order to achieve this, we can introduce two indicator functions, one for using a diffusion autoencoder and another for using a standard VAE. Let’s define $\mathbb{I}_{\text{diff}^{(m)}}$ as the indicator function for using a diffusion autoencoder for modality $m$, and $\mathbb{I}_{\text{vae}^{(m)}}$ as the indicator function for using a standard VAE for modality $m$. Then, we can express the objective as shown in equation 5.
$\begin{split}\sum_{A\in\mathcal{S}}\omega_{A}\Bigg{\{}&\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x}_{A})}\Bigg{[}\sum_{i}^{M}\mathbb{I}_{\text{vae}^{(i)}}\log p_{\theta}(\mathbf{x}_{i}|\mathbf{z})+\mathbb{I}_{\text{diff}^{(i)}}\mathbb{E}_{\mathbf{x_{i},\epsilon_{t}}}\Big{[}\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{i_{t}},\mathbf{z},t)\right\|^{2}\Big{]}\Bigg{]}\\ &-D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x}_{A})\|p(\mathbf{z}))\Bigg{\}}\end{split}$ (5)
The above formulation may not still be a valid lower bound on the log-likelihood, but it provides a seamless way to integrate the diffusion decoder into the multimodal VAE setup. We choose the product of experts as the posterior where one subset is sampled from the powerset of modalities. In addition to this, to increase the alignment between representations of different modalities, we add a KL-regularization term between the posterior and the individual encoder outputs. The final objective is shown in equation 6 where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters set to $1e^{-5}$ because higher values caused instability during training. We call our proposed model, which combines diffusion autoencoders and variational autoencoders for multimodal generative modeling, Multimodal Diffusion Decoder Variational Autoencoder (MDDVAE).
$\begin{split}\sum_{A\in\mathcal{S}}\omega_{A}\Bigg{\{}&\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x}_{A})}\Bigg{[}\sum_{i}^{M}\mathbb{I}_{\text{vae}^{(i)}}\log p_{\theta}(\mathbf{x}_{i}|\mathbf{z})+\mathbb{I}_{\text{diff}^{(i)}}\mathbb{E}_{\mathbf{x_{i},\epsilon_{t}}}\Big{[}\left\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{i_{t}},\mathbf{z},t)\right\|^{2}\Big{]}\Bigg{]}\\ &-\lambda_{1}D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x}_{A})\|p(\mathbf{z}))-\lambda_{2}\sum_{i}^{M}\Big{[}D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x}_{A})\|q(\mathbf{z}|\mathbf{x}_{i}))\Big{]}\Bigg{\}}\end{split}$ (6)
3.3 Conditional Sampling
To perform conditional sampling (i.e. generate missing modalities given some modalities) utilizing our model, we can leverage the factorized nature of the posterior distribution. We first sample from the joint posterior distribution conditioned on the available modalities and then use the sampled latent variable to generate samples using the respective generative model for each modality. The posterior will be the product of the observed modalities $q_{\phi}(\mathbf{z}|\mathbf{x}_{o})=\prod_{i\in{o}}q_{\phi}(\mathbf{z}_{i}|\mathbf{x}_{i})$ where the subscript $o$ denotes observed modalities. Then, we can generate an unobserved modality ${u}$ from the posterior $q_{\phi}(\mathbf{z}|\mathbf{x}_{o})$ by feeding it to the appropriate decoder.
3.4 Unconditional Sampling
Unconditional generation happens when no specific modality is provided as input. In this scenario, the goal is to produce samples that exhibit coherence across all modalities without relying on any initial conditions. Previous multimodal VAEs perform unconditional generation by first sampling from a $\mathcal{N}(0,\mathbf{I})$ and then feeding the sample to each decoder. Doing that may produce suboptimal results as there is a gap between the posterior and the prior (Aneja et al., 2021). To get better samples, we train an auxiliary score-based diffusion model to sample from the posterior. Figure  2 demonstrates how the score model facilitates unconditional sampling. This auxiliary score model is trained using the diffusion objective but on the latent space $\mathbf{z}$ to generate samples from the posterior. The objective is $\mathbb{E}_{t,\mathbf{z}_{0},\boldsymbol{\epsilon}}[\left\|\boldsymbol{\epsilon}-\epsilon_{\theta}(\mathbf{z}_{t},t)\right\|^{2}$ where $\mathbf{z_{0}}$ is sampled from $q_{\phi}(\mathbf{z}|\mathbf{x})=\prod_{i}q_{\phi}(\mathbf{z}_{i}|\mathbf{x}_{i})$ where $\mathbf{x}_{i}$ consists of each modality in $\mathbf{x}$. The main objective of training this auxiliary model is to get high-quality samples from the posterior instead of using $\mathcal{N}(0,\mathbf{I})$ as the starting sample for unconditional generation. To generate samples unconditionally, we first sample $\mathbf{z}$ from the auxiliary score model and use the respective generative model $p_{\theta}(\mathbf{x}_{i}|\mathbf{z})$ of each modality to get the final samples.
4 Experiments
We perform experiments on two datasets. The first one is from Wah et al. (2011), which contains images of birds with text captions describing the birds. We use actual images of the birds instead of ResNET features, unlike previous multimodal VAEs that struggle with image data (Shi et al., 2019). The second dataset is a high-dimensional CelebAMask-HQ  (Lee et al., 2020) dataset that consists of images, masks, and attributes of celebrities expressed in the three modalities. We compare different multimodal VAE baselines consisting of MVTCAE by Hwang et al. (2021), MoPoE by Sutter et al. (2021), and MMVAE+ Palumbo et al. (2023). MMVAE+ is not included in the CUB dataset experiment as we use pre-trained VAE for the text captions, which will require changing the structure of the text VAE because of the modality-specific and shared representations in MMVAE+. Nonetheless, we added the results from their paper for the text-to-image results, which can be referenced from their work (Palumbo et al., 2023). We also add additional baselines from  Wesego and Rooshenas (2023); Bounoua et al. (2024) that use a latent score-based model to learn the joint latent space discussed in the next sections.
4.1 Caltech Birds (CUB)
The CUB dataset consists of two modalities: image and text describing the image (caption). In order to get meaningful text outputs, we initialized the text VAE from the pre-trained weight of  Li et al. (2020) that uses a BERT encoder and GPT-2 decoder for all models. The images are resized to 64x64. We used a latent size of $768$ for both the image and text. The encoder architecture for the image modality is similar to that of  Daniel and Tamar (2021) and the diffusion decoder uses a UNET architecture. The models were trained for 500 epochs using the Adam optimizer (Kingma and Ba, 2015). The baseline models are trained using different $\beta$ values listed in the appendix, and the best model was selected by using the average of the conditional and unconditional FID. The unconditional auxiliary score model also uses a 1D UNET architecture and accepts input similar to the size of the latent dimension. DDIM sampling was used for 50 steps to generate samples for the image modality and for unconditional sampling (Song et al., 2021). More details about the training and inference are in the Appendix section.
a small sized bird that has multiple tones of grey all over
this bird has a pointed bill with a yellow breast and belly
Figure 4: Conditional generation given a text using MDDVAE
Table 1: CUB FID Result
Txt-to-Img Unc
MDDVAE 35.2($\pm$0.57) 37.5($\pm$0.09)
MoPoE 290.6($\pm$0.5) 199.6($\pm$0.75)
MVTCAE 176.3($\pm$0.05) 168.9($\pm$0.2)
MLD 62.6 63.4
MMVAE+ 164.9 -
We use FID, which is a widely adopted metric for evaluating the quality of images, as the evaluation metric to measure conditionally and unconditionally generated samples (Heusel et al., 2017). In addition to that, we use Clip-Score to measure the similarity of the image-text pairs even if it’s not the best metric for this dataset  (Hessel et al., 2021; Bounoua et al., 2024). It’s clear that a good multimodal VAE model will have coherent unconditional samples where the image and text represent the same concept expressed in its respective modality. In addition, conditionally generated images should follow the description in the text, and conditionally generated text should follow the given image.
bird has gray body feathers white breast feather and black beak
a small with a white belly brown breast and crown and a yellow throat
Figure 6: Unconditional generation using MDDVAE
We present the results in Table  1 and  2. In addition to the implemented baselines, we add results from previous works of MLD by  Bounoua et al. (2024) that uses latent diffusion model and MMVAE+ on text-to-image FID. The FID scores evaluation demonstrates that our model, MDDVAE, outperforms the baselines, generating high-quality images that exhibit greater coherence and alignment with the text modality. In contrast, the baselines struggle to generate meaningful images that are consistent with the textual information, supporting the findings of Daunhawer et al. (2022) regarding the challenges faced by existing multimodal VAEs in difficult tasks. Table  2 also shows that MDDVAE generates more coherent conditional outputs that follow the given modality. We also added qualitative results from our model, which is the only multimodal VAE model generating good images in this dataset, are shown in figure  4 and  6 for conditional and unconditional generation respectively.
Table 2: CUB Clip Score Result
Img-to-Txt Txt-to-Img Unc
MDDVAE 28.1($\pm$0.001) 28.52($\pm$0.02) 28.12($\pm$0.005)
MoPoE 26.8($\pm$0.03) 22.0($\pm$0.03) 24.8($\pm$0.05)
MVTCAE 27.1($\pm$0.03) 26.6($\pm$0.004) 28.08($\pm$0.02)
4.2 CelebAMask-HQ
The CelebAMask-HQ dataset provides multimodal representations of visual characteristics related to human faces. In this dataset, the images, masks, and attributes can be viewed as distinct yet complementary modalities that capture various aspects of an individual’s appearance. We resize the images and masks to a resolution of 128 by 128 pixels. We combine multiple masks of parts of the face into a single 0/1 mask similar to  Wesego and Rooshenas (2023). Specifically, all the provided masks in the CelebAMask-HQ dataset, except for the skin mask, are drawn on top of each other, resulting in a single composite mask. This composite mask encodes the presence or absence of various facial features. For the attribute modality, we follow the preprocessing approach described by Wu and Goodman (2018); Wesego and Rooshenas (2023). Out of the 40 existing attributes in the CelebAMask-HQ dataset, we selectively choose 18 attributes that are most relevant and informative for representing the visual characteristics of a person’s face. We used a latent size of $256$. Similar to the previous dataset, the encoder and decoder architecture for the image modality are inspired by  Daniel and Tamar (2021) except for the diffusion decoder, which uses a UNET architecture. The unconditional auxiliary score-based model also works on the $256$ dimension latent size resized to 16x16 utilizing a UNET architecture. Sampling is done by using the DDIM sampling technique for 50 steps (Song et al., 2021). More details about the experimental setup are located in the appendix section. The baselines here are MoPoE, MVTCAE, MMVAE+ plus SBM-VAE-C and SBM-RAE-C by  Wesego and Rooshenas (2023) and MLD by  Bounoua et al. (2024), which use unimodal variational autoencoders (VAEs) or autoencoders (AEs) for each modality and fuse them using a score-based model for learning the joint distribution.
(a) MDDVAE
(b) MVTCAE
(c) MMVAE+
(d) MoPoE
Figure 8: Conditional generated images for different models given the same masks and attribute for all the models
Table 3: CelebAMask-HQ Result
Attribute Image Mask
GIVEN Both Img Both Mask Attr Unc Both Img
F1 F1 FID FID FID FID F1 F1
MDDVAE 0.76($\pm$0.001) 0.76($\pm$0.001) 32.4($\pm$0.08) 32.4($\pm$0.06) 28.3($\pm$0.27) 32.4($\pm$0.25) 0.90($\pm$0.001) 0.90($\pm$0.001)
MoPoE 0.68($\pm$0.002) 0.71($\pm$0.004) 114.9($\pm$0.32) 101.1($\pm$0.16) 186.7($\pm$0.28) 164.8($\pm$0.62) 0.85($\pm$0.002) 0.92($\pm$0.001)
MVTCAE 0.71($\pm$0.001) 0.69($\pm$0.004) 94($\pm$0.45) 84.2($\pm$0.32) 87.2($\pm$0.08 162.2($\pm$1.08) 0.89($\pm$0.001) 0.89($\pm$0.003)
MMVAE+ 0.64($\pm$0.003) 0.61($\pm$0.002) 133($\pm$14.28) 97.3($\pm$0.04) 153($\pm$0.49) 103.7($\pm$0.61) 0.82($\pm$0.03) 0.89($\pm$0.002)
SBM-RAE-C 0.66($\pm$0.03) 0.64($\pm$0.004) 83.6($\pm$0.05) 82.8($\pm$0.05) 83.1($\pm$0.14) 84.2($\pm$0.05) 0.83($\pm$0.004) 0.82($\pm$0.004)
SBM-VAE-C 0.69($\pm$0.005) 0.66($\pm$0.001) 82.4($\pm$0.1) 81.7($\pm$0.29) 76.3($\pm$0.7) 79.1($\pm$±0.3) 0.84($\pm$0.02) 0.84($\pm$0.001)
MLD 0.71($\pm$0.005) 0.67($\pm$0.006) 81.7($\pm$0.25) 82.4($\pm$0.15) 80.29($\pm$0.6) 82.8($\pm$±0.08) 0.86($\pm$0.001) 0.86($\pm$0.001)
Supervised 0.79($\pm$0.001) 0.94($\pm$0.001)
To assess the performance of the models in generating high-quality samples across different modalities, we employ two quantitative evaluation metrics. For the image modality, we utilize the Fréchet Inception Distance (FID) score. On the other hand, for the mask and attribute modalities, we employ the sample-average F1 score as the evaluation metric. The F1 score provides a comprehensive measure of the model’s ability to accurately generate binary predictions in multi-label classifications.
We present our results in table  3 which shows the performance of the models in both conditional and unconditional settings. MDDVAE generates high-quality images, achieving the best FID scores among all models in image generation. Not only that, but having a better representation model in the image modality also helps the performance of the model in the other modalities where MDDVAE outperforms the baselines in attribute prediction, even if the attribute modality architectures are the same for all models. Our models also have very close results to supervised prediction models that are trained to predict the attribute or mask directly from the image modality taken from  Wesego and Rooshenas (2023). Unconditional results also show that the baselines do not generate good-quality images unconditionally. The effect of the auxiliary score-based model can be seen in this case, where MDDVAE can leverage it to generate coherent outputs along all the modalities. In general, the results indicate that our proposed MDDVAE model exhibits superior performance in both conditional and unconditional generation tasks, consistently producing higher-quality and more coherent outputs compared to the baselines.
Finally, in addition to quantitative results, we show qualitative results in Figure  8. The figure shows conditionally generated images for MDDVAE and the baselines. MDDVAE generates higher-quality images that are more coherent with the given conditions.
4.3 Ablation: Effect of the Auxiliary Score Model
Figure 10: Unconditional generation using MDDVAE when using no auxiliary score model (left) and when using auxiliary score model (right)
In this section, we will look at the importance of using the auxiliary score-based model. By default, multimodal VAEs sample from $\mathcal{N}(0,\mathbf{I})$ to perform unconditional generation. The problem with this is that there could be considerable differences in distribution between the prior and the posterior, which will lead to low-quality samples. Previous unimodal VAE works improved sample quality by training an energy-based model on the latent space Aneja et al. (2021); Pang et al. (2020). The energy-based model will serve as a prior instead of the standard normal distribution and provide samples closer to the posterior, thereby improving generated outputs.  Preechakul et al. (2022) also used a score-based model in the unimodal diffusion autoencoder to generate samples unconditionally. In light of these, we train an auxiliary score-based model to generate prior samples for unconditional multi-modal generation. The auxiliary score-based model will transform the standard normal sample to a sample from the posterior $q_{\phi}(\mathbf{z}|\mathbf{x})$. Figure  10 shows the effect of using the auxiliary score model. The left side of the figure shows outputs without using the auxiliary score model. Those outputs are inferior compared with the right-side ones, which use the auxiliary score model and generate high-quality outputs.
5 Conclusion and Discussion
In this paper, we introduce the Multimodal Diffusion Decoder Variational Autoencoder (MDDVAE), which enhances the capabilities of traditional multimodal VAEs by incorporating diffusion decoders, leading to superior performance. Additionally, we implement an auxiliary score-based model to bridge the gap between the prior and posterior distributions, achieving high-quality and coherent results even in unconditional scenarios. A notable limitation of MDDVAE is its higher computational demand during sampling due to the iterative diffusion process, despite the use of DDIM sampling to reduce the sampling time. Nonetheless, our model demonstrates significant performance improvements across multiple challenging datasets, advancing the field of multimodal VAEs and positioning them as a preferred choice for multimodal applications. Future work will focus on optimizing the multimodal model to reduce computational overhead. Our contributions provide a robust framework for future research and development in multimodal VAE generative modeling.